Grade 12th passAlgebra

Saima wants to put lace on the edge of a circular table cover of a diameter of 1.5 m. Find the length of the lace required and also find its cost if one meter of the lace costs ₹ 15. (Take π = 3.14)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that,
Diameter of the circular table = 1.5 m
We know that, radius (r) = d/2
= 1.5/2
= 0.75 m
Then,
Circumference of the circle = 2πr
= 2 × 3.14 × 0.75
= 4.71 m
So, the length of lace = 4.71 m
Cost of 1 m lace =₹15 [given]
Cost of 4.71 m lace =₹15 × 4.71
=₹70.65

Thanks